Order: Dopamine so meglkg min w Available: 800mg in 200ML DSW pt. weighs 132 lbs. What is the flow rate?
Answer #1

1 pound = 0.454 kg

132 pounds = 132 * 0.454 = 59.93 kg

Dose = 50 mcg * 59.93 = 2996.5 mcg = 3 mg

Dose is 3 mg/min

800 mg in 200 ml

3 mg in = 200/800 * 3 = 0.75 ml/min